OpenStudy (allieeslabae):

help with 2 problems?

OpenStudy (mjdennis):

Sure. #24 first. You know cosine = (x/r) for the 'standard position', right? How do they define secant and cosecant?

OpenStudy (allieeslabae):

For 24. I did the work already and got B.

OpenStudy (allieeslabae):

\[\csc \theta =\frac{ 5 }{ 4 }, \sec \theta=\frac{ 5 }{ 3 }\]

OpenStudy (allieeslabae):

@mjdennis am I right, or no?

OpenStudy (mjdennis):

You are right on B. Oh. What have you done on 25? I'll get caught up in a minute

OpenStudy (allieeslabae):

MMhm nothing so far I dont know where to start. Guide me?

OpenStudy (mjdennis):

OK, it's like a jigsaw. We know they are 'sine functions' right? Even if it it cosine, there is an identity that turns cosine into sine, so we'll start with the most generic sine function: f(theta) = A1*sin(b1*theta + phase1) g(theta) = A2*sin(b2*theta + phase2) start with f(0) = 0 and g(0) = 0 to find phase1 and phase2

OpenStudy (allieeslabae):

If this doesnt sound dumb, may I ask how do we do that.. lol I suck at these.

OpenStudy (mjdennis):

 But that is OK, we can skip that part here.